OpenStudy (anonymous):

A rectangular box is twice as long as it is wide. The height of the box is 3 feet less than the width. If the box is x feet wide, what polynomial represents its volume in cubic feet? A)4x-3 B)2x^3+6^2 C)2x^2-6x D)2x^3-6x^2

OpenStudy (anonymous):

^ is powered aka exponent

OpenStudy (kropot72):

Let the width = w Then the height = w - 3 And the length = 2w Can you use the above definitions to form an equation for the volume?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Do you want to post your equation?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x-3 2(2x-3) ??? did i set it up right

OpenStudy (kropot72):

The volume of the box is: V = w * 2w * (w-3) = 2w^2(w - 3) = 2w^3 - 6w^2 Substituting x for w: V = 2x^3 - 6x^2 So the correct choice is D

OpenStudy (anonymous):

thank u

OpenStudy (kropot72):

You're welcome :)